OpenStudy (anonymous):

Calculus question, medal for help and explanation. Much appreciated I will post below. thanks

OpenStudy (anonymous):

\[ f(t)=\int\limits_{1}^{t^2}(\sqrt{7+u^6}du/u \] and \[ F(x)=\int\limits_{1}^{x}f(t)dt, \] Find F''(2). Thanks and id like to understand it as well

OpenStudy (anonymous):

I also dont have the answer so il try my answer and if someone could confirm thatd be good to. so F'(x)=f(x) so \[f(x)=\int\limits_{1}^{x^2}(\sqrt{7+u^6du/u}\] so F''(x)= d/dx of that integral, which would be sqrt(7+x^2)^6/x^2 times 2x and evaluated at x=2

OpenStudy (anonymous):

OK, first let's make clearance about your problem. I would like to understand the f(t) \[F(x) = \int\limits_{1}^{x}f(t)dt\] So F'(x) = f(t) with t from 1--> x. And F''(x) = f'(t) But f'(t) can be written as\[f' (t) = \frac{\sqrt{7+u^6}}{u}\] with u from 1 --> t^2 That's almost done. Now just do some simplification to the final result. Step 1: substitute u=1 and u=t^2 in order to calculate f'(t) \[f \prime (t) = \frac{\sqrt{7+t^{12}}}{t^2} - \sqrt{8}\] Step 2: substitute t =1 and t=x in order to calculate F''(x) = f'(t), below I already simplify the expression for you. \[F \prime \prime (x) = \frac{\sqrt{7+x^{12}}}{x^2} - \sqrt{8}\] Step 3: F''(2) is ? \[F \prime \prime (2) = \frac{\sqrt{4103}}{4}-\sqrt{8} \approx 13.185\]

OpenStudy (anonymous):

Where did your - sqrt come from

OpenStudy (anonymous):

You mean -sqrt(8) ? This was taken from definition of definite integrals \[F(x) = \int\limits_{a}^{b} f(t)dt = F(b) - F(a)\]

OpenStudy (anonymous):

Okay and do you not multiple by 2x because of the x^2 and chain rule

OpenStudy (anonymous):

So do you agree from the steps: F''(x) = f'(t) with t = 1-->x ? With substitute t = x then the chain rule cannot apply.

OpenStudy (anonymous):

More clearly, let me explain by : \[F \prime (x) = f(t)\] Take derivative of d/dx for both sides \[\frac{d}{dx}F'(x) = \frac{d}{dx}f(t) = f'(t) . \frac{dt}{dx}\] Here is the chain rule application, but because t = x at boundary then dt/dx = 1

OpenStudy (anonymous):

Alright makes sense, thanks